### v1

# General Exploratory Bonus for Optimistic Exploration in RLHF

Abstract
Optimistic exploration is central to improving sample efficiency in reinforcement learning with human feedback, yet existing exploratory bonus methods to incentivize exploration often fail to realize optimism. We provide a theoretical analysis showing that current formulations, under KL or $\alpha$-divergence regularization, unintentionally bias exploration toward high-probability regions of the reference model, thereby reinforcing conservative behavior instead of promoting discovery of uncertain regions. To address this pitfall, we introduce the General Exploratory Bonus (GEB), a novel theoretical framework that provably satisfies the optimism principle. GEB counteracts divergence-induced bias via reference-dependent reward regulation and unifies prior heuristic bonuses as special cases, while extending naturally across the full $\alpha$-divergence family. Empirically, GEB consistently outperforms baselines on alignment tasks across multiple divergence settings and large language model backbones. These results demonstrate that GEB offers both a principled and practical solution for optimistic exploration in RLHF. Code is available here.
1 Introduction
Despite the acknowledged significance of online exploration for reinforcement learning with human feedback (RLHF) (Xu et al., 2024; Tang et al., 2024; Tajwar et al., 2024), there remains a paucity of theoretical frameworks governing how to explore. As shown in Fig. 1 (1, top), standard online RLHF algorithms (Guo et al., 2024; Pang et al., 2024; Ahmadian et al., 2024) generally rely on passive exploration, i.e., the stochasticity of the policy itself to generate responses, with no mechanism to incentivize novelty or diversity. As a result, this approach can be notoriously sample-inefficient. When the optimal behavior resides in low-probability regions, passive exploration is unlikely to discover it, leading to policies that remain trapped around local optima.
To address this, some works (Das et al., 2024; Ji et al., 2024; Mehta et al., 2023; Qi et al., 2025; Liu et al., 2024) have attempted to devise sample-efficient algorithms, inspired by the principle optimism in the face of uncertainty. As illustrated in Fig. 1 (2, top), the principle aims to generate responses with high epistemic uncertainty, thus encouraging data collection in unexplored regions for further training. To operationalize the principle, recent research (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025) encourages exploration by adding exploratory bonuses to the reward modeling, which is practically optimizable for large language models. These methods intend to artificially inflate rewards in underexplored regions, nudging the policy toward more informative data collection.
Unfortunately, our theoretical analysis in Section 3 reveals a fundamental pitfall: under the common KL-regularized RLHF, the existing theoretical framework of exploratory bonuses fails to satisfy optimism. In particular, we prove that existing bonus formulations can undesirably drive the policy $\pi$ toward the reference policy $\pi_{\textrm{ref}}$ due to the divergence regulation in the exploratory bonus, and the induced bonus actually biases exploration toward high-probability regions of the reference model. As illustrated in Fig. 1 (II, bottom), the bonus disproportionately amplifies rewards for regions already well-covered by $\pi_{\textrm{ref}}$, thereby reinforcing conservative behavior rather than driving exploration into uncertain regions. This failure is not confined to KL-divergence; we further extend our analysis to the more general $\alpha$-divergence family and prove that the same collapse persists across a wide range of divergence-regularized objectives. Thus, while existing approaches appear to encourage exploration, they in fact undermine the very principle of optimism they aim to realize.
Motivated by these failures, we propose a new framework, General Exploratory Bonus (GEB), which theoretically unifies existing approaches while provably satisfying optimism (Section 4). GEB corrects the failure modes of prior approaches by directly introducing a reference-dependent regulation into the reward. This adjustment offsets the undesired conservatism induced by divergence regularization, allowing the exploratory bonus to satisfy optimism—it increases the probability of responses rarely sampled to pursue potentially more preferred answers, as shown in Fig. 1 (III, bottom). Importantly, GEB provides a unified formulation: prior heuristic exploratory bonuses can be reinterpreted as special cases, and the framework naturally extends to the full class of $\alpha$-divergences. Beyond correcting the theoretical shortcomings, GEB remains practically implementable—it can be seamlessly integrated into the standard iterative RLHF loop without additional sampling cost.
We validate GEB on a large-scale alignment task across different divergences and model backbones. Empirically, GEB consistently yields stronger alignment compared to its counterpart of passive exploration. For example, the three GEB variants that we consider generally outperform the iterative f-DPO (Xiong et al., 2024a) across different divergence regulations, while the most performant variant surpasses several existing optimistic exploration methods that incorporate exploratory bonuses (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025). By analyzing the distribution of sampled responses, we validate that GEB can successfully encourage sampling in the region of small $\pi_{\textrm{ref}}$, thereby effectively achieving optimistic exploration.
We summarize our main contributions:
  1. 1.
    We formally prove that the existing theoretical framework of exploratory bonuses under KL and $\alpha$-divergence regularization fails to achieve optimistic exploration.
  2. 2.
    We introduce General Exploratory Bonus (GEB), a novel theoretical framework of optimistic exploration for RLHF that provably satisfies the optimism principle and unifies prior heuristic bonuses.
  3. 3.
    We empirically validate GEB on LLM alignment tasks, showing improved performance and broad applicability across multiple divergence families.
Figure 1: The upper part compares passive exploration and optimistic exploration. Optimistic exploration stimulates the trajectories $\tau$ of small $\pi_{\textrm{ref}}$ (seldom visited/uncertain). While passive exploration sticks to the high-$\pi_{\textrm{ref}}$ region, failing to approach global optima. The lower part contrasts the effect of the exploration bonus term in optimistic reward modeling between prior works and our GEB. Prior works often emphasize rewards in frequently visited regions, which constrains exploration within certain areas. In contrast, our GEB amplifies rewards in seldom-visited regions, thereby encouraging further sampling in uncertain areas and successfully achieving optimistic exploration.
2 Preliminaries
Iterative online RLHF. The effectiveness of iterative online RLHF (Xiong et al., 2024b; Dong et al., 2024) has been validated in many real-world systems such as Claude (Bai et al., 2022) and LLaMA-series (Touvron et al., 2023; Grattafiori et al., 2024). The algorithm proceeds in $T$ rounds, with each round having two steps: (1) the $\pi_{t}$ is learned with the current dataset $\mathcal{D}_{t}$, and then samples $x\sim\rho,\ (y_{1},y_{2})\sim\pi_{t}(\cdot|x)$; (2) Human evaluator annotate the preference of $(x,y_{1},y_{2})\rightarrow(x,y^{w},y^{l})$ to form $\mathcal{D}_{t+1}$, where the prompt $x$ is sampled from an independent distribution $\rho$, response $y_{1},y_{2}$ are two response sampled from the policy of the $t$-th iteration $\pi_{t}$. When computing $\pi_{t}$ with dataset $D_{t}$ in the step (1) of each iteration, a reward function $r_{t}(x,y)$ is first learned from a collected human preference data $\mathcal{D}_{t}=\{(x,y^{w},y^{l})\}$, where $y^{w},y^{l}$ denote the preferred and dispreferred response to $x$, respectively. Reward modeling typically follows the Bradley-Terry objective (Bradley & Terry, 1952):
$r_{t}=\arg\min_{r}\mathcal{L}_{BT}(\mathcal{D}_{t},r)=\arg\min_{r}\mathbb{E}_{(x,y^{w},y^{l})\sim\mathcal{D}_{t}}-\log[\sigma(r(x,y^{w})-r(x,y^{l}))],$ (1)
where $\sigma$ denotes the sigmoid function. Given the learned reward function $r_{t}$, the policy $\pi_{t}$ is then updated to maximize the expected reward, often with a KL-regularization as follows
$\pi_{t}=\arg\max_{\pi}\mathcal{J}_{\beta,\text{KL}}(\pi,r_{t})=\arg\max_{\pi}\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}r_{t}(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}}),$ (2)
where $\beta>0$ is a hyperparameter, and $\pi_{\textrm{ref}}$ is the reference model.
Sample inefficiency of iterative online RLHF. Online sampling for standard online RLHF algorithms is carried out passively, relying solely on the inherent randomness of the LLM policy. However, if the policy places a small probability mass on the optimal action, passive exploration may fail to ever explore this action. Theoretical analyses (Muldrew et al., 2024; Dong et al., 2024) and empirical evidence (Chen et al., 2025; Cen et al., 2025) present that the passive approach fails to sufficiently explore the prompt-response space. Particularly, Xie et al. (2024) demonstrate that the sample complexity can be exponential in $1/\beta$ for passive exploration, which is unacceptable in the small-$\beta$ regime. Therefore, several works (Xie et al., 2024; Cen et al., 2025; Zhang et al., 2024a) propose exploratory bonuses to implement optimistic exploration for efficient sampling. However, in the next section, we will show that prior formulations cannot provably achieve optimism.
3 Exploratory Bonus and How It Can Fail
In this section, we will first provide the iterative online RLHF formulation with an exploratory bonus (Section 3.1). We then theoretically prove that the existing formulation can fail to achieve optimistic exploration under both KL-constrained RLHF (Section 3.2) and a more general $\alpha$-divergence-regularized RLHF (Section 3.3), motivating our proposed method in Section 4.
3.1 Exploratory Bonus
To improve the sample efficiency of iterative online RLHF, recent works (Zhang et al., 2024a; Cen et al., 2025) introduce exploratory bonuses, which try to encourage optimistic exploration. These approaches modify the standard loop by adding an exploratory bonus term $\mathcal{L}_{\text{bonus}}$. Specifically, in the $t$-th iteration, the reward model $r_{t}$ and policy $\pi_{t}$ are optimized by
$r_{t}$ $=\arg\min_{r}\Big[\mathcal{L}_{BT}(\mathcal{D}_{t},r)-\kappa\mathcal{L}_{\text{bonus}}(r)\Big],$ (3)
$\pi_{t}=\arg\max_{\pi}$ $\mathcal{J}_{\beta,\text{KL}}(\pi,r_{t})=\arg\max_{\pi}\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}r_{t}(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}}),$ (4)
where $\kappa>0$ is a hyperparameter. By Eq. 3, the reward model $r_{t}$ should not only fit the observed data in $\mathcal{D}_{t}$, but also learn to maximize the bonus term $\mathcal{L}_{\text{bonus}}(r)$.
To achieve optimistic exploration, the bonus term is expected to stimulate probability increase more prominently in unexplored areas. Formally, we have the following definition:
Definition 3.1 (Optimism condition for exploration bonus)
When a reward model $r$ and a policy $\pi$ is computed with Eq. 3 and Eq. 4, the exploratory bonus $\mathcal{L}_{\text{bonus}}(r)$ achieves optimism, if
$\frac{\partial^{2}\mathcal{L}_{\text{bonus}}(r)}{\partial\pi(y|x)\partial\pi_{s}(y|x)}<0,$ (5)
where $\pi_{s}(y|x)$ is an ideal sampling distribution for response at the current iteration.
To interpret the optimism condition of the exploration bonus term in Definition. 3.1, we consider the policy-reparameterized reward model $r(x,y):=r(\pi)$, which can be derived from the closed-form solution of Eq. 2 as $\pi(y|x)=\frac{\exp(r(x,y)/\beta)}{Z(x)}$, where $Z(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}}\exp(r(x,y)/\beta)$ is a normalization function. This yields the reward model expressed via the policy (Rafailov et al., 2024):
$r(\pi)=\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}+\log Z(x).$
Implication.
Eq. 5 requires the gradient $\partial\mathcal{L}_{\text{bonus}}(r(\pi))/\partial\pi(y|x)$ to be negatively correlated with the sampling probability $\pi_{s}$. In other words, responses rarely sampled under $\pi_{s}$ (i.e., uncertain or underexplored outputs) should receive a larger ascending of the policy distribution $\pi$, i.e., larger $\partial\mathcal{L}_{\text{bonus}}(r(\pi))/\partial\pi(y|x)$. In practice, $\pi_{s}$ can be substituted by the reference model or intermediate checkpoints. We adopt the commonly used $\pi_{\textrm{ref}}$ as $\pi_{s}$ in our following demonstration.
3.2 Failure Under KL-constrained RLHF
Previous works, including Zhang et al. (2024a) and Cen et al. (2025), formulate the exploratory bonus with $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,KL}(\pi,r)$. In this case, optimizing exploratory bonus in Eq. 3 becomes a min-max bi-level objective as $\min_{r}-\kappa\max_{\pi}[\mathbb{E}_{x,y\sim\pi}r(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}})]$. Intuitively, they intend to make $r$ not only fit the observed data by $\mathcal{L}_{BT}$, but also have a larger reward in unobserved regions by maximizing the $\max_{\pi}\mathbb{E}_{x,y\sim\pi}r(x,y)$ in $\mathcal{L}_{\text{bonus}}(r)$. Here, we theoretically show that such formulations can suffer from optimism failures under KL-regularized RLHF.
Lemma 3.1 (Optimism failure under KL-divergence.)
Let $r_{1}=\arg\min_{r}\mathcal{L}_{BT}(\mathcal{D},r)$ be a reward model trained with the vanilla BT loss, and let $r_{2}=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\max_{\pi}\mathcal{J}_{\beta,\text{KL}}(\pi,r)]$ be a reward model trained with an additional exploratory bonus. If the policy is optimized via Eq. 4, then $r_{1}$ and $r_{2}$ yield the same set of policies.
Proof  First, the inner maximization of the bonus term admits a closed-form solution, $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)e^{\frac{r(x,y)}{\beta}}/Z(x)$ where $Z(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}e^{\frac{r(x,y)}{\beta}}$ is a normalization term. Substituting this solution reduces the bi-level training objective of $r_{2}$ to a single-level form:
$r_{2}=\arg\min_{r}\big[\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathbb{E}_{x\sim\rho}\beta\log\mathbb{E}_{y\sim\pi_{\textrm{ref}}}e^{\frac{r(x,y)}{\beta}}].$ (6)
As shown in Rafailov et al. (2024), the log-ratio $\beta\log\pi_{\theta}(y|x)-\beta\log\pi_{\textrm{ref}}(y|x)$ represents the same class of the original reward function $r$ through Eq. 4, thus the reward $r_{1},r_{2}$ can be reparameterized by the log-ratio. Plugging this into Eq. 6 yields
$\arg\min_{\pi}\mathcal{L}_{dpo}(\mathcal{D},\pi)-\kappa\mathbb{E}_{x\sim\rho}\beta\log\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}.$ (7)
Since the second term equals 0, the reparameterized Eq. 6 is exactly the vanilla DPO loss, which is the same as the reparameterized training objective of $r_{1}$. Thus, the exploratory bonus in the reward training objective has no effect on the final policy set. $\square$
The lemma proves that incorporating the exploratory bonus $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,\text{KL}}(\pi,r)$ into the reward training objective fails to induce the policy model to sample from low-$\pi_{\textrm{ref}}(y|x)$ regions, i.e., unexplored responses. In other words, the bonus term is ineffective at inducing optimism. We next extend the result beyond KL regularization to a more general class of $\alpha$-divergence family.
3.3 Generalization to $\alpha$-divergence-constrained RLHF
In this subsection, we theoretically show that the failure of optimism can broadly extend to the $\alpha$-divergence class. Many common divergences, such as reverse KL-divergence, Hellinger distance, and forward KL-divergence, are special cases of $\alpha$-divergence.
Definition 3.2 ($\alpha$-divergence class)
An $\alpha$-divergence is a certain type of function $D(p|q)=\int f(\frac{\mathrm{d}p}{\mathrm{d}q})\mathrm{d}q$ that measures the difference between two probability distributions $p$ and $q$, where
$f(x)=\frac{x^{\alpha}-\alpha x-(1-\alpha)}{\alpha(1-\alpha)},$
and $\alpha$ is a hyperparameter typically with $0\leq\alpha\leq 1$.
Lemma 3.2 (Optimism failure under $\alpha$-divergence.)
Consider objective $\mathcal{J}_{\beta,f}(\pi,r)=\mathbb{E}_{x\sim\rho,y\sim\pi(y|x)}r(x,y)+\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}f(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, where $f$ belongs to $\alpha$-divergence class. If a reward is trained with $r=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathcal{L}_{\text{bonus}}]$ and a policy $\pi$ is updated by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ with $\mathcal{L}_{\text{bonus}}=\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$, the gradient of the bonus satisfies $\frac{\partial^{2}\mathcal{L}_{\text{bonus}}(r(\pi))}{\partial\pi\partial\pi_{\textrm{ref}}}\geq 0$, which means $\mathcal{L}_{\text{bonus}}$ encourage trajectories with large $\pi_{\textrm{ref}}$ more strongly, in contradiction to the optimism principle (Definition 3.1).
Proof  For a RL objective $\mathcal{J}_{\beta,f}(\pi,r)$, the relation between the optimal policy $\pi^{*}_{f}$ and the reward $r$ can be formulated as follows,
$\pi_{f}^{*}=\frac{1}{Z(x)}\pi_{\textrm{ref}}(y|x)(f^{\prime})^{-1}(r/\beta),\quad r(x,y)=\beta f^{\prime}(\frac{\pi^{*}(y|x)}{\pi_{\textrm{ref}}(y|x)}Z(x)),$ (8)
where $Z(x)$ is a normalization term and $(f^{\prime})^{-1}$ is the inverse function of $f^{\prime}$. The bi-level objective can be similarly transformed to a single level one by canceling the inner maximization $\max_{\pi}$ by Eq. 8. The single-level objective can be written as $r_{t}=\arg\min_{r}\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{1}{Z(x)}(f^{\prime})^{-1}(\frac{r(x,y)}{\beta})\cdot r(x,y)-\beta f(\frac{1}{Z(x)}(f^{\prime})^{-1}(\frac{r(x,y)}{\beta})$. Since the policy is computed by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$, the reward can be reparameterized by the policy with Eq. 8, which fortunately cancels $Z(x)$. Then the optimistic reward-modeling objective can be reparameterized as
$\arg\min_{\pi}\mathcal{L}_{dpo}(\mathcal{D},\pi)-\kappa\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\Big[\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}f^{\prime}(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})-\beta f(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})\Big].$ (9)
Since for $\alpha$-divergence, $f(u)=\frac{u^{\alpha}-\alpha u-(1-\alpha)}{\alpha(\alpha-1)}$, the partial derivative of Eq. 9 is $(\frac{\pi_{\textrm{ref}}}{\pi})^{1-\alpha}$, which induces positively correlated gradients w.r.t. $\pi$ and $\pi_{\textrm{ref}}$ when $0\leq\alpha<1$, and is a constant when $\alpha=1$, hence contradictory to the optimism defined in Definition 3.1. $\square$
Table 1: Realized exploratory bonus under different divergence classes when $\mathcal{L}_{\text{bonus}}(r)=\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$.
$f$ exploratory bonus
reverse KL constant
forward KL $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}$
Hellinger distance $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\sqrt{\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}}$
According to Lemma. 3.2, we enumerate several exploratory bonus under different $\alpha$-divergence in Table 1. The listed bonuses are simplified by removing constant coefficients and bias. The listed exploratory bonuses generally force the policy model to maximize the possibility of trajectories sampled by the reference model, not the underexplored ones. We further prove that it actually drives $\pi$ to collapse toward $\pi_{\textrm{ref}}$ and that the failure extends beyond $\alpha$-divergence to other $f$-divergences.
Theorem 3.3 (Optimism failure beyond $\alpha$-divergence.)
When $f$ belongs to $f$-divergence, and the reward function is obtained by $\hat{r}=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D}_{t},r)-\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)]$ and the policy is updated by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r_{t})$, the bonus term $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ induces the policy model $\pi$ to coincide with $\pi_{\textrm{ref}}$ when $xf^{\prime\prime}(x)$ is a monotone function.
The detailed proofs are in Appendix B.1. The monotone increase of $xf^{\prime\prime}(x)$ can be satisfied by a broader divergence class besides $\alpha$-divergence, including JS-divergence and Pearson $\chi^{2}$.
Intuitive understanding.
Optimization of exploratory bonus in Eq. 3 is a min-max bi-level objective as $\min_{r}-\kappa\max_{\pi}[\mathbb{E}_{x,y\sim\pi}r(x,y)-\beta\mathbb{D}_{\text{KL}}(\pi\|\pi_{\textrm{ref}})]$. Due to the inner maximization $\max_{\pi}$, the divergence constraint implicitly makes $\pi$ close to $\pi_{\textrm{ref}}$ to avoid the divergence penalty. Considering the outer minimization $\min_{r}$, $r$ is forced to provide large rewards on region of high $\pi$ to maximize the reward expectation. Their combination implicitly makes $r$ focus more on region of large $\pi_{\textrm{ref}}$. Since samples with large $\pi_{\textrm{ref}}$ is easily rolled out from scratch, previous exploratory bonuses merely concentrate sampling on regions that are already easy to explore, contradictory to the optimism principle, which requires encouraging responses $y$ rarely sampled by the reference model.
4 General Exploratory Bonus with Optimism Principle
Motivated by the failure of the existing exploratory bonus, we now propose a novel framework, General Exploratory Bonus (GEB), and prove that it achieves optimism. We further show that prior exploratory bonuses—and broader variants—emerge as special cases of our formulation.
Formulation of a novel exploratory bonus.
As shown in Section 3, the failure of existing bonuses arises because the divergence constraints in $\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ force $\pi$ to align with $\pi_{\textrm{ref}}$, biasing exploration toward high-$\pi_{\text{ref}}$ regions. To achieve optimistic exploration, the optimal $\pi$ must instead counteract this regularization and move away from $\pi_{\textrm{ref}}$. Our key idea is to introduce an additional reference-dependent regulation into the reward, which offsets the influence of divergence regularization. The resulting exploratory bonus takes the form $\kappa\max_{\pi}J_{\beta,f}(\pi,R(r,\pi_{\textrm{ref}}))$. Note that the formulation of $R(\cdot,\cdot)$ can be diverse. Since the reward $R(r,\pi_{\textrm{ref}})$ is now explicitly dependent on $\pi_{\textrm{ref}}$, the inner optimal policy of the bi-level problem is $\frac{1}{Z_{R}(x)}\pi_{\textrm{ref}}(f^{\prime})^{-1}(\frac{R(r,\pi_{\textrm{ref}})}{\beta})$, where $Z_{R}(x)$ is a normalization term. Unlike previous cases, the optimal policy is no longer guaranteed to be positively correlated with $\pi_{\textrm{ref}}$, enabling the policy to deviate from the reference distribution.
As in Lemma 3.2, we can substitute the inner $\pi$ by the closed-form solution, and then utilize the reward reparameterization of $\alpha$-divergence (Wang et al., 2024) as $r=f^{\prime}(\pi/\pi_{\textrm{ref}})$ for a divergence instance $f$ to obtain the reparameterized exploratory bonus as
$\mathcal{L}_{\text{bonus}}(r(\pi))=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(\cdot|x)}\Big[\frac{u}{Z_{R}(x)}f^{\prime}(u)-f(\frac{u}{Z_{R}(x)})\Big],$ (10)
where $u=(f^{\prime})^{-1}(R(\pi_{\textrm{ref}}(y|x),\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})/\beta)$ and $Z_{R}(x)=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}u$. Since the domain of divergence class is generally $(0,+\infty)$ while there are no additional constraints on the formulation of $R$, $u$ can be flexibly formulated with $\pi$ and $\pi_{\textrm{ref}}$ unless $u\geq 0$.
Equivalence to a practical objective.
In our proposed exploratory bonus, the normalization term $Z_{R}(x)$ in Eq. 10 cannot be canceled. Fortunately, we prove the following lemma in Appendix E to show the equivalence between the two training objectives, one with and the other without $Z_{R}(x)$, which helps transform the objective to a succinct formulation for analyses and practical use.
Lemma 4.1
Denote two objectives as $h(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}uf^{\prime}(u)-f(u)$ and $\hat{h}(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{u}{Z(x)}f^{\prime}(u)-f(\frac{u}{Z(x)})$ where $u$ is a function with $\pi$ and $\pi_{\textrm{ref}}$. When $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z(x)})]/[Z(x)uf^{\prime\prime}(u)]=\Lambda(x)$ is constant in $y$ and $\Lambda(x)>0$, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induce the same class of policies.
GEB successfully achieves optimism.
Building on Lemma 4.1, we now prove that our proposed framework indeed achieves the optimism requirement.
Theorem 4.2
For each iteration of online RLHF, if the policy is updated by $\pi=\arg\max_{\pi}J_{\beta,f}(\pi,r)$ while its reward is trained with $r=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D},r)-\kappa\mathcal{L}_{\text{bonus}}]$. When $f$ belongs to $\alpha$-divergence class, and $\mathcal{L}_{\text{bonus}}=\max_{\pi}J_{\beta,f}(\pi,R(r,\pi_{\textrm{ref}}))$, denote $u(\pi,\pi_{\textrm{ref}})=(f^{\prime})^{-1}\Big(R\Big((f^{\prime})^{-1}(\frac{\pi}{\pi_{\textrm{ref}}}),\pi_{\textrm{ref}}\Big)/\beta\Big)$, Then, $\frac{\partial^{2}\mathcal{L}_{\text{bonus}}}{\partial\pi\partial\pi_{\textrm{ref}}}\leq 0$ if $\forall(x,y);\frac{\partial u}{\partial\pi}+\pi_{\textrm{ref}}\frac{\partial^{2}u}{\partial\pi\partial\pi_{\textrm{ref}}}+\frac{(\alpha-1)\pi_{\textrm{ref}}}{u}\frac{\partial u}{\partial\pi}\frac{\partial u}{\partial\pi_{\textrm{ref}}}<0$ and $u>\alpha$, where $\alpha$ is a hyperparameter that defines $\alpha$-divergence (Definition 3.2).
Proof  First, substituting the optimal solution of the inner maximization and utilizing reward reparameterization, we obtain the training objective as in Eq. 10. By Lemma 4.1, this can be equivalently expressed as
$\mathcal{L}_{\text{bonus}}=\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\Big[uf^{\prime}(u)-f(u)\Big].$ (11)
For $\alpha$-divergences, the conditions in Lemma 4.1 are satisfied. Since $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z})=u^{\alpha-1}(Z^{1-\alpha}-\alpha)/(1-\alpha)$ and $uf^{\prime\prime}(u)=u^{\alpha-1}$, the fraction $\Lambda(x)=(Z^{1-\alpha}-\alpha)/Z(1-\alpha)$ in Lemma 4.1 is independent of $y$. Since $u>\alpha$, $Z_{R}=\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}u>0$, thus $\Lambda(x)>0$ is also satisfied. Finally, the mixed second-order derivative of Eq. 11 is computed as
$\frac{\partial^{2}\mathcal{L}_{\text{bonus}}}{\partial\pi\partial\pi_{\textrm{ref}}}=\beta\mathbb{E}_{x\sim\rho}\sum_{y}u^{\alpha-1}(\frac{\partial u}{\partial\pi}+\pi_{\textrm{ref}}\frac{\partial^{2}u}{\partial\pi\partial\pi_{\textrm{ref}}}+\frac{(\alpha-1)\pi_{\textrm{ref}}}{u}\frac{\partial u}{\partial\pi}\frac{\partial u}{\partial\pi_{\textrm{ref}}})<0,$ (12)
which achieves the optimism defined in Definition 3.1. $\square$
Table 2: General exploratory bonus under different divergence classes and design of $u$. Note all $u>\alpha$ when $0<\pi<1$. The presented bonus is simplified by removing constant coefficients or biases.
$f$ $u$ $\mathcal{L}_{\text{bonus}}$ $1+\alpha-\pi$ $1/\pi$ $\mathrm{arctanh}(1-\pi)+\alpha$
reverse KL $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}-\pi(y|x)$ $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{1}{\pi(y|x)}$ $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\mathrm{arctanh}(1-\pi(y|x))$
forward KL $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\log(1-\pi(y|x))$ $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}-\log\pi(y|x)$ $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\log\mathrm{arctanh}(1-\pi(y|x))$
Hellinger Distance $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\sqrt{1.5-\pi(y|x)}$ $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{1}{\sqrt{\pi(y|x)}}$ $\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\sqrt{\mathrm{arctanh}(1-\pi(y|x))+0.5}$
In our formulation, $u$ can be flexibly defined in terms of $\pi,\pi_{\textrm{ref}}$ as long as it satisfies the derivative condition in Theorem 4.2 and $u>\alpha$. This flexibility highlights the extensibility of our framework. In particular, when $u$ depends only on $\pi$, any function with $u>\alpha$ and negative correlation with $\pi$ qualifies. In Table 2, we list several such choices of $u$, along with their corresponding reparameterized exploratory bonus under three different $\alpha$-divergences.
From a practical standpoint, since $\mathcal{L}_{\text{bonus}}$ is expressed as an expectation over $\pi_{\textrm{ref}}$, it does not require additional sampling and can be seamlessly integrated into iterative online RLHF. To avoid unintended decreases in the likelihood of preferred responses, however, we follow Chen et al. (2025) and restrict computation of the bonus on rejected responses to ensure that the probability of preferred responses continues to increase.
Prior exploratory bonuses are encompassed within GEB.
Although we have shown that existing theoretical formulations of $\mathcal{L}_{\text{bonus}}$ fail to guarantee optimism, many practical implementations have nevertheless been effective through various approximations and adaptations. These approximations and adaptations are generally inextensible beyond the reverse KL divergence (detailed in Appendix B.2). In this subsection, we show that these practical implementations can be naturally subsumed into our GEB framework, and even broader objectives can be reinterpreted as instances of optimistic exploration. For example, Zhang et al. (2024a) and Xie et al. (2024) finally implement their exploratory bonus as $\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(y|x)}\log\pi(y|x)$, which belongs to GEB when $u(\pi)=-\log\pi+1$ and $f$ is KL-divergence. Similarly, Cen et al. (2025) implement the exploratory bonus as $\kappa\mathbb{E}_{x\sim\rho,y\sim\pi_{\text{cal}}(\cdot|x)}\log\frac{\pi}{\pi_{\textrm{ref}}}$ where $\pi_{\text{cal}}$ is a fixed calibration distribution. This also falls under GEB by setting $u=-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log\frac{\pi}{\pi_{\textrm{ref}}}-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log{\pi_{\textrm{ref}}}+1$ and $f$ is KL-divergence. The corresponding reparameterized exploratory bonus reduces to $\mathbb{E}_{x\sim\rho,y\sim\pi_{cal}}-\log\frac{\pi}{\pi_{\textrm{ref}}}+C(x)$, where $C(x)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}[-\frac{\pi_{\text{cal}}}{\pi_{\textrm{ref}}}\log{\pi_{\textrm{ref}}}+1]$. Interestingly, even objectives not explicitly designed for exploration can be reinterpreted through our GEB framework. For instance, Chen et al. (2025) augment the DPO loss with an additional term $\kappa\mathbb{E}_{x,y\sim\pi_{ref}}\sigma(-\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, which was originally introduced to control sample complexity. In our framework, this corresponds to optimistic exploration with $u=-\sigma(-\beta\log\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})+1$.
5 Experiments
Table 3: In-domain evaluation on different exploration bonuses. Boldface and underline indicate the best and the second-best results, respectively. $\pi, GEB$-$1/\pi$, and GEB-$\mathrm{arctanh}(\pi-1)$ corresponds to $1+\alpha-\pi$, $1/\pi$, and $\mathrm{arctanh}(1-\pi)+\alpha$ as in Table 2.
KL ($\alpha$=1) Hel. ($\alpha$=0.5) f-KL ($\alpha$=0) Avg.
WR AvgR WR AvgR WR AvgR WR AvgR
Mistral-Instruct-v0.3
f-DPO 78.42 0.7480 72.69 0.6536 51.11 0.5918 67.40 0.6645
SELM 77.56 0.7530 - - - - - -
XPO 79.71 0.7492 - - - - - -
VPO 78.57 0.7426 - - - - - -
FEB 78.42 0.7480 71.54 0.6525 47.53 0.5928 65.83 0.6644
GEB-$\pi$ 81.00 0.7542 75.48 0.6641 51.68 0.5976 69.39 0.6720
GEB-$1/\pi$ 80.00 0.7554 73.97 0.6541 52.26 0.6051 68.74 0.6715
GEB-$\mathrm{arctanh}(\pi-1)$ 79.71 0.7559 75.69 0.6614 52.76 0.5989 69.39 0.6721
LLaMA-3-8B-SFT
f-DPO 73.11 0.8050 71.11 0.7859 67.38 0.7579 70.53 0.7829
SELM 74.19 0.8126 - - - - - -
XPO 72.40 0.8119 - - - - - -
VPO 71.61 0.7971 - - - - - -
FEB 73.11 0.8050 68.17 0.7591 67.95 0.7611 69.74 0.7751
GEB-$\pi$ 74.34 0.8156 71.68 0.7840 67.67 0.7681 71.23 0.7892
GEB-$1/\pi$ 74.76 0.8102 72.25 0.7859 68.17 0.7591 71.73 0.7851
GEB-$\mathrm{arctanh}(\pi-1)$ 74.98 0.8080 73.26 0.7877 68.89 0.7569 72.38 0.7842
Table 4: Out-of-domain evaluation on different exploration bonuses with LLaMA-3-8B-SFT. Boldface and underline indicate the best and the second-best results, respectively. $\pi, GEB$-$1/\pi$, and GEB-$\mathrm{arctanh}(\pi-1)$ corresponds to $1+\alpha-\pi$, $1/\pi$, and $\mathrm{arctanh}(1-\pi)+\alpha$ as in Table 2.
KL($\alpha$=1) Hel.($\alpha$=0.5) f-KL($\alpha$=0) Avg.
Alpaca Math Alpaca Math Alpaca Math Alpaca Math
f-DPO 25.72 67.6 24.73 69.0 17.80 69.2 22.75 68.6
FEB 25.72 67.6 23.75 68.6 19.62 68.6 23.03 68.3
GEB-$\pi$ 28.27 69.2 25.87 69.6 20.05 71.6 24.73 70.1
GEB-$1/\pi$ 26.10 68.4 25.28 70.2 19.80 69.2 23.73 69.3
GEB-$\mathrm{arctanh}(\pi-1)$ 24.90 71.0 25.96 67.6 19.62 69.2 23.49 69.3
5.1 Experimental Settings
Following prior works (Zhang et al., 2024a; Xie et al., 2024; Chen et al., 2025), we adopt the same iterative online algorithm as in Algorithm 1 with three iterations, aiming to isolate the effects of different exploration bonuses. We adopt two LLM backbones: Llama-3-8B-SFT (Dong et al., 2024) following prior works, and Mistral-Instruct-v0.3 (Jiang et al., 2023). The training prompt set is RLHFlow-UltraFeedback (Dong et al., 2024) as in previous works. URM-LLaMa-3.1-8B (Lou et al., 2024) serves as the preference oracle. We evaluate the outcome policies on both in-domain and out-of-domain test sets. Specifically, for the in-domain test, we use a held-out test set from UltraFeedback (Cui et al., 2024), and sample 64 times per prompt with the outcome policy to compare the average reward and win-rates against the base model. We use length-controlled AlpacaEval2 benchmark (Dubois et al., 2024) with GPT-4 as a judge for out-of-domain alignment test, and MATH-500 (Lightman et al., 2023) to evaluate out-of-domain reasoning ability.
Baselines.
We adopt f-DPO (Wang et al., 2024), which extends DPO to the f-divergence class, as the primary baseline. We further compare GEB with three optimistic-exploration methods that incorporate exploratory bonuses—SELM (Zhang et al., 2024a), XPO (Xie et al., 2024), and VPO (Cen et al., 2025). Since the approximations or adaptations in their implementations do not extend beyond the KL divergence, we report their results only under KL. In contrast, we introduce a new baseline, Failed Exploratory Bonus (FEB), which removes these approximations or adaptations, i.e., Eq. 7.
5.2 Results & analyses
GEB delivers robust improvements across different loss designs, divergence classes, and language model backbones.
The experimental results are shown in Table 3. Across both backbones, GEB generally outperforms f-DPO and FEB. Under the KL-divergence, GEB displays better or at least on-par performance compared to prior exploratory-bonus methods. Notably, the win-rate increases over 1.82% and 0.94% under the KL-divergence, over 2.36% and 1.29% under the Hellinger Distance, compared with their f-DPO counterpart. GPT-4 evaluation on the Alpaca benchmark also shows consistent performance gains on out-of-domain alignment task. While GEB maintains on par, or usually better results in MATH, showing less performance degradation beyond alignment, known as alignment tax (Noukhovitch et al., 2023; Lin et al., 2024).
Figure 2: Comparison of $\log\pi_{\textrm{ref}}$ of sampled response in the last iteration between the general exploratory bonuses and vanilla iterative DPO. $\pi, GEB$-$1/\pi$, and GEB-$\mathrm{arctanh}(\pi-1)$ corresponds to $1+\alpha-\pi$, $1/\pi$, and $\mathrm{arctanh}(1-\pi)+\alpha$ as in Table 2
Figure 3: Experiments with different $\kappa$. The three graphs are under KL divergence, Hellinger Distance, and forward KL divergence from left to right, respectively. The p, f, tanh in the legends correspond to $1+\alpha-\pi$, $1/\pi$, $\mathrm{arctanh}(1-\pi)+\alpha$ in Table 2 respectively.
Table 5: Dist-n of the sampled corpus in the last iteration under the KL divergence.
dist-1 dist-2 dist-3 dist-4
f-DPO 0.0189 0.2700 0.6349 0.8418
GEB-$\pi$ 0.0192 0.2694 0.6323 0.8420
GEB-$1/\pi$ 0.0191 0.2738 0.6401 0.8448
GEB-$\mathrm{arctanh}(\pi-1)$ 0.0192 0.2730 0.6391 0.8447
GEB effectively encourages exploration in small $\pi_{\textrm{ref}}$ region, yielding more diverse sampling.
In Figure 2, we visualize the distribution of $\log\pi_{\textrm{ref}}$ for sampled responses in the last iteration under the KL divergence. When trained with the GEB, the policy model consistently samples more trajectories with a smaller $\pi_{\textrm{ref}}$ compared to the policy trained by f-DPO loss. This validates our motivation that GEB can encourage sampling trajectories of small $\pi_{\text{ref}}$ for optimistic exploration. In Table 5, we further calculate the distinct-n ($n=1,2,3,4$) for the sampled responses in the last iterations under the KL divergence, which measures the diversity of a corpus. GEB generally has higher diversity scores, validating that GEB incentivizes qualitatively more diverse samples.
The choice of $\kappa$.
Since the formulation of $u$ in Eq. 10 is flexible, the scale of the GEB term can differ substantially across designs, hence the absolute value of the bonus is less informative. Instead, we examine the relative ratio of the bonus term to the vanilla RL loss $|\kappa\mathcal{L}_{bonus}|/|\mathcal{L}_{RL}|$, which provides a more consistent basis for comparison and offers better practical guidance for tuning $\kappa$ across diverse settings. As shown in Fig. 3, performance remains stable when the ratio lies within a suitable range (1e-2 to 1e-6 in our case). However, if the ratio is too large, it impedes optimization of the RL objective and degrades performance; if too small, the exploration incentive in uncertain regions diminishes and performance reverts to the vanilla baseline.
6 Conclusion
While recent work proposes exploratory bonuses to operationalize the “optimism in the face of uncertainty” principle, our work shows that the existing theoretical frameworks of exploratory bonuses fail under KL and $\alpha$-divergence regularization. To address prior theoretical pitfalls, we introduce General Exploratory Bonus (GEB), a novel theoretical framework for sample-efficient RLHF. Our approach provably satisfies the optimism principle and unifies prior heuristic bonuses. We empirically validate GEB on LLM alignment tasks with diverse bonus designs and LLM backbones, showing improved performance and broad applicability across multiple divergence families.
Acknowledgement
The authors would like to thank Froilan Choi and Shawn Im for their valuable comments on the manuscript. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23-1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon.
Reproducibility statement
We have included all implementation details, hyperparameters, and training procedures in the paper and appendix. Our code and scripts for reproducing the experiments are publicly available through the GitHub repository.
Ethics statement
This work studies reinforcement learning from human feedback (RLHF) using only publicly available or synthetic data, without new human subject collection. Here, by providing a rigorous theoretical framework with strong empirical evidence, we pursue a high standard of scientific excellence. We also take into account inclusiveness to make all our visualizations accessible to the unprivileged group of people, by producing figures distinguished by light, shade, and marker. While RLHF has the potential to amplify biases or harmful behaviors if misused, our work is intended solely to advance safe and responsible research, and we encourage its application in alignment with ethical standards.
Appendix A Related Works
Alignment & RLHF.
Alignment (Hendrycks et al., 2021; Leike et al., 2018; Zhang et al., 2024b; 2025; Yeh et al., 2025) aims to ensure AI systems act in accordance with human values, preferences, and goals; and it has become a critical field in AI research. To steer language models to match human preferences, reinforcement Learning from Human Feedback (RLHF) (Stiennon et al., 2020; Ouyang et al., 2022) acheives great success and has become the standard alignment pipeline. However, its computational complexity has motivated a family of Direct Preference Optimization (DPO) (Rafailov et al., 2024; Azar et al., 2024; Ethayarajh et al., 2024) that forgo explicit reward modeling. Despite their efficiency, recent researchers (Tajwar et al., 2024; Xu et al., 2024; Xiong et al., 2024b) reemphasize the significance of online sampling.
Optimistic exploration of RLHF.
To address the computational overheads of passive exploration in RLHF, which samples trajectories just based on randomness, some existing attempts have been devoted to sample-efficient RL algorithms. Most of works (Das et al., 2024; Ji et al., 2024; Dwaracherla et al., 2024; Mehta et al., 2023; Muldrew et al., 2024) adhere to the principle of optimism, proposing specialized prompt or response selection strategies to emphasize uncertain samples. While some research (Liu et al., 2024; Lou et al., 2024) propose uncertainty-aware reward models with epistemic neural networks or bootstrap ensembles, these methods introduce additional cost. Some research also addresses the sample efficiency with different theoretical foundations, such as information theory (Qi et al., 2025), preference-incentive exploration (Chen et al., 2025). Notably, several works (Zhang et al., 2024a; Xie et al., 2024; Cen et al., 2025) introduce different exploratory bonuses, which can implement optimism toward uncertainty without additional computes. However, they only focus on KL-divergence and their theoretical framework cannot result in real optimism as shown in Section 3.2.
Efficient RL for LLM.
Beyond optimistic exploration, some research proposes fine-grained signals for RL learning. For instance, several research propose segment-level (Kong et al., 2025; Yin et al., 2025) or token-level (Zeng et al., 2024; Li et al., 2024) reward function for alignment or text control. Notably, for reasoning tasks, process reward model (Lightman et al., 2024; Li & Li, 2025; Yuan et al., 2024) which provides step-wise feedback for solutions has shown promise effectiveness. On the other hand, recent research (Cheng et al., 2025; Zheng et al., 2025; Wu et al., 2025) on LLM reasoning reveal that high-entropy tokens guide the model toward diverse reasoning paths. Training with only high-entropy tokens are more beneficial for reasoning performance (Wang et al., 2025). While our approach is highly extensible, we believe the orthogonal methods can be further incorporated with our general exploratory bonus.
Appendix B Optimism Failure of previous works
B.1 Extension beyond $\alpha$-divergence
The following theorem formally proves that the exploratory bonus $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ cannot encourage optimism for a more general divergence class.
Theorem 3.3
When $f$ belongs to $f$-divergence, and the reward function is obtained by $\hat{r}=\arg\min_{r}[\mathcal{L}_{BT}(\mathcal{D}_{t},r)-\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)]$ and the policy is updated by $\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r_{t})$, the bonus term $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ induces the policy model $\pi$ to coincide with $\pi_{\textrm{ref}}$ when $xf^{\prime\prime}(x)$ is a monotone function.
Proof  By Lemma 3.2, we reparameterize the bonus term for optimistic reward-modeling to Eq. 8. Denote $h(u)=uf^{\prime}(u)-f(u)$. For a fixed prompt $x$, we formulate the training process as a constrained problem as follows,
$\arg\max_{\pi}\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}h(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})\quad s.t.\quad\sum_{y}\pi(y|x)=1\quad\textrm{and}\quad\forall y,\pi(y|x)>0.$ (13)
Then we can apply the Lagrange multiplier as
$\mathcal{L}=\mathbb{E}_{y\sim\pi_{\textrm{ref}}}h(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})-\mu\Big(\sum_{y}\pi(y|x)-1\Big)-\sum_{y}\eta(y)\pi(y|x),$ (14)
where $\mu,\eta$ are the dual variables. Then we utilize the Karush-Kuhn-Tucker (KKT) conditions for the given optimization problem. The complementary slackness requires that $\forall y,\eta(y)\pi(y|x)=0$. The stationary condition requires
$\frac{\partial\mathcal{L}}{\partial\pi(y|x)}=h^{\prime}(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})-\mu-\eta(y)=0.$ (15)
We denote $S_{y}=\{y|\pi(y|x)>0\}$, we have $\forall y\in S_{y},\eta(y)=0$. Since $h^{\prime}(u)=uf^{\prime\prime}(u)>0$ due to the convexity of $f(\cdot)$, we can obtain $\forall y\in S_{y},\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}$ is a constant. Then applying the normalisation constraint, $\mathbb{E}_{y\sim\pi_{\textrm{ref}}(\cdot|x)}\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)}=1$. Hence, the unique interior optimum is $\pi^{*}(y|x)=\pi_{\textrm{ref}}(y|x)$. $\square$
The theorem implies that the reparameterized exploratory bonus attains its maximum only when $\pi$ and $\pi_{\textrm{ref}}$ coincide. The condition that $xf^{\prime\prime}(x)$ is a monotone function is satisfied by $\alpha$-divergence and beyond, e.g. Pearson $\chi^{2}$. Hence, the exploratory bonus $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ in the reward training objective generally contradicts the optimism, since it cannot encourage trajectories with small initialized possibility.
B.2 Prior adaptions & approximations cannot generalize
Though the theoretical framework of the prior exploratory bonus fails, their empirically implemented loss remains effective through different adaptations and approximations. However, in this subsection, we show that these adaptations and approximations are inextensible beyond the KL-divergence class.
Zhang et al. (2024a) adapt the formulation of $\mathcal{J}_{\beta,f}(\pi,r)$ in $\kappa\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r)$ as
$\mathcal{J}^{\prime}_{\beta,f}(\pi,r)=E_{x,y\sim\pi,y^{\prime}\sim\pi_{\textrm{ref}}}[r(x,y)-r(x,y^{\prime})]-\beta D_{KL}(\pi|\pi_{\textrm{ref}}),$ (16)
which adds a bias in the reward expectation term. Under KL-divergence, the original $\mathcal{J}_{\beta,f}(\pi,r)$ will be zero after re-parameterization as shown in Lemma 3.1, thus the sole reparameterized bias term will remain as $\mathcal{E}_{y^{\prime}\sim\pi_{\textrm{ref}}}\log\pi(y^{\prime}|x)$. Since $\mathcal{J}_{\beta,f}(\pi,r)$ cannot be reparameterized to zero except for the KL-divergence, this adaptation cannot generalize then.
In the derivations of Cen et al. (2025), it utilizes an ideal distribution $\pi_{cal}$ which should satisfy $\mathbb{E}_{y\sim\pi_{cal}}r(x,y)=0$. Since $\pi_{cal}$ is practically unobtainable, it uses the rejected responses to approximate $\mathbb{E}_{\pi_{cal}}$, which does not satisfy the predefined condition of $\pi_{cal}$, thus not rigorously coherent to the theory. While the regret decomposition of Xie et al. (2024) relies on the logarithm form of the KL-divergence, it is thus inextensible to a broader divergence class.
In contrast, our general exploratory bonus can seamlessly incorporate an iterative online RLHF algorithm and can naturally extend to the entire $\alpha$-divergence class. All prior bonuses mentioned above can be encompassed by our theoretical framework.
Appendix C Regret Bound
In derivations, we utilize the theoretical tools in Zhang et al. (2024a); Xie et al. (2024); Cen et al. (2025). First, we make some standard statistical assumptions following Cen et al. (2025).
Assumption C.1
For a reward function $r$, and a random function $R(\cdot)$, and any trajectory $\tau$, we have $R_{max}<r(\tau),R(r(\tau))<R_{max}$, where $R_{max}$ is a constant.
This is an assumption generally made for theoretical analyses of RLHF. Note that $R_{max}$ is measurable and controllable in practice. Then we introduce the assumption of the reward class proposed in Cen et al. (2025), which offers a regularization mechanism to incorporate additional policy preferences in the subsequent derivations.
Assumption C.2
We assume that $r^{*}\in\mathcal{R}$, where
$\mathcal{R}=\{r:\mathbb{E}_{x\sim\rho,\tau\sim\pi_{cal}(\cdot|x)}r(x,y)=0\},$ (17)
where $\rho$ is the prompt distribution and $\pi_{cal}$ is a fixed calibration distribution independent of the algorithm.
We also introduce the preference generalized eluder coefficient proposed in Zhang et al. (2024a), an extension of the generalized eluder coefficient (Zhong et al., 2022), which connects prediction error and in-sample estimation error.
Definition C.1
Let $f_{r}(x,y,y^{\prime})=r_{t}(x,y)-r^{*}(x,y)$. For a reward function class $\mathcal{R}$, we define the preference generalized eluder coefficient as the smallest $d_{\mathrm{PGEC}}$ as
$\sum_{t=1}^{T}$ $\mathbb{E}_{x\sim\rho,y\sim\pi^{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]$
$\leq\sqrt{d_{\mathrm{PGEC}}\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho,y\sim\widetilde{\pi}_{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2}}+4\sqrt{d_{\mathrm{PGEC}}T},$ (18)
With the above assumptions and the theoretical tool, we can have the following regret boundary.
Theorem C.1
Let $\mathcal{J}_{\beta,f}(\pi,r)=\mathbb{E}_{x\sim\rho,y\sim\pi(\cdot|x)}r(x,y)-\beta\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}(\cdot|x)}f(\frac{\pi(y|x)}{\pi_{\textrm{ref}}(y|x)})$, when the hyperparameter of the loss Eq. 3 $\kappa=\sqrt{\frac{\log(T|\mathcal{R}|\delta^{-1})}{(\gamma d_{\mathrm{PGEC}}T)}}(32R_{max}e^{4R_{max}})^{-1}$, with probability at least $1-\delta$, the regret can be bounded as follows,
$\sum_{t=1}^{T}\mathcal{J}_{\beta,f}(\pi^{*},r^{*})-\mathcal{J}_{\beta,f}(\pi^{t},r^{*})\leq\mathcal{O}(R_{max}e^{4R_{max}}T\sqrt{d_{\mathrm{PGEC}}\gamma\log(|\mathcal{R}|\delta^{-1})}),$ (19)
where $\gamma=\sup_{x,y}\frac{\pi}{\pi_{cal}}$, $r^{*}$ and $\pi^{*}$ are ground-truth reward function and corresponding optimal policy with $\pi^{*}=\arg\max_{\pi}\mathcal{J}_{\beta,f}(\pi,r^{*})$.
Proof  First, we can decompose the regret function as in Cen et al. (2025) as follows,
$\underbrace{\sum_{t=1}^{T}[\mathcal{J}_{\beta,f}(\pi^{*},r^{*})-\mathcal{J}_{\beta,f}(\pi^{t},r^{t})]}_{\textrm{Term 1}}+\underbrace{\sum_{t=1}^{T}[\mathcal{J}_{\beta,f}(\pi^{t},r^{t})-\mathcal{J}_{\beta,f}(\pi^{t},r^{*})]}_{\textrm{Term 2}}.$ (20)
Then, we will bound term 1 and term 2 individually and combine them at last.
Bound term 1.
First, we connect the term 1 with $\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r_{t}))-\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r^{*}))$. When $\pi^{*}$ is the optimal $\pi$ for $\max_{\pi}J_{\beta,f}(\pi,r^{*})$, we have
$\textrm{Term 1}\leq J_{\beta,f}(\pi^{*},r^{*})-J_{\beta,f}(\pi^{*},r_{t})\leq\sup_{x,y}\frac{\pi^{*}}{\pi_{t}}\mathbb{E}_{x\sim\rho,y\sim\pi_{t}}(r^{*}-r^{t}).$ (21)
Similarly, we can obtain its lower bound as $2R_{max}$. Then, we have
$\textrm{Term 1}\leq\sum_{t=1}^{T}\Big[\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r_{t}))$ $-\max_{\pi}\mathcal{J}_{\beta,f}(\pi,R(r^{*}))\Big]+4R_{max}T.$ (22)
Bound term 2.
First, we utilize the preference generalized eluder coefficient to connect the prediction error to in-sample error.
$\textrm{Term 2}=\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho,y\sim\pi^{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]$ (23)
$\leq\frac{\eta Td_{\mathrm{PGEC}}}{4}+4\sqrt{{d_{\mathrm{PGEC}}}T}+\frac{1}{\eta}\sum_{t=1}^{T}\mathbb{E}_{x\sim\rho,y\sim\widetilde{\pi}_{t}(\cdot|x),y^{\prime}\sim\pi_{cal}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2},$ (24)
where the first equality uses the property of the reward class in Assumption C.2, and the inequality follows Definition C.1 with Cauchy–Schwarz inequality. Then, we bound the squared in-sample error as follows.
$\mathbb{E}_{x\sim\rho,\tau\sim\widetilde{\pi}_{t}}$ $[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2}\leq\gamma\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}}[f_{r_{t}}(x,y,y^{\prime})-f_{r^{*}}(x,y,y^{\prime})]^{2}$ (25)
$\leq\gamma(32R_{max}e^{4R_{max}})^{2}\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}}[\sigma(f_{r_{t}}(x,y,y^{\prime}))-\sigma(f_{r^{*}}(x,y,y^{\prime}))]^{2}$ (26)
$\leq 8\gamma(32R_{max}e^{4R_{max}})^{2}\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}(\cdot|x)}D_{H}^{2}(P_{r_{t}}(\cdot|\tau,\tau^{\prime})\|P_{r^{*}}(\cdot|\tau,\tau^{\prime})),$ (27)
where $\gamma=\sup_{x,y}\frac{\widetilde{\pi}_{t}}{\pi_{cal}}$, and the second inequality utilizes the Lemma C.8 in Xie et al. (2024), the third inequality uses $(x-y)^{2}<4(x+y)(\sqrt{x}-\sqrt{y})$. Refer to the Lemma C.6 in Xie et al. (2024), we have
$\sum_{i<t}\mathbb{E}_{x\sim\rho,\tau,\tau^{\prime}\sim\widetilde{\pi}_{t}(\cdot|x)}D_{H}^{2}(P_{r_{t}}(\cdot|\tau,\tau^{\prime})\|P_{r^{*}}(\cdot|\tau,\tau^{\prime}))\leq L_{BT}^{(t)}(r_{t})-L_{BT}^{(t)}(r^{*})+2\log(|\mathcal{R}|\delta^{-1})$ (28)
where $L_{BT}^{(t)}(r)=\sum_{i<t}\mathbb{E}_{y,y^{\prime}\sim\mathcal{D_{t}}}-\log\sigma(f_{r}(x,y,y^{\prime}))$ is the vanilla BT loss for reward modeling. Finally, the term 2 can be bounded by
$\textrm{Term 2}\leq 4\sqrt{{d_{\mathrm{PGEC}}}T}$ $+\frac{\eta Td_{\mathrm{PGEC}}}{4}+$
$\frac{8\gamma}{\eta}(32R_{max}e^{4R_{max}})^{2}(L_{BT}^{(t)}(r_{t})-L_{BT}^{(t)}(r^{*})+2T\log(|\mathcal{R}|\delta^{-1})).$ (29)
Bound the regret.
Since the $r_{t}$ is optimized by $L_{BT}^{(t)}(r_{t})-\sum_{i=1}^{T}\kappa\max_{\pi}\mathcal{J}_{\beta,f}(R(r_{t}),\pi)$, we have $r_{t}=\arg\min_{r\in\mathcal{R}}L_{BT}^{(t)}(r_{t})-\sum_{i=1}^{T}\kappa\max_{\pi}\mathcal{J}_{\beta,f}(R(r_{t}),\pi)$ Therefore, when $\eta=4\sqrt{\frac{2\gamma\log(T|\mathcal{R}|\delta^{-1})}{Td_{\mathrm{PGEC}}}}(32R_{max}e^{4R_{max}})$ and $\kappa=\frac{\eta}{4\gamma}(32R_{max}e^{4R_{max}})^{-2}$, the regret can be bounded by
Regret $\leq 4\sqrt{{d_{\mathrm{PGEC}}}T}+\sqrt{2^{3}\gamma d_{\mathrm{PGEC}}\log(|\mathcal{R}|\delta^{-1})}(32R_{max}e^{4R_{max}}T)+4R_{max}T.$ (30)
$\square$
Appendix D Experiments
D.1 Implementation Details
Algorithm.
Following prior works Zhang et al. (2024a); Xie et al. (2024); Chen et al. (2025), we adopt the same algorithmic backbone for empirical validation to explicitly show the effect of different exploratory bonuses in the loss function. This algorithm bypasses the reward modeling in each iteration through reward reparameterization, known as iterative DPO (Dong et al., 2024). Previous works further reparameterize the bonus term to incorporate the algorithm. Since iterative DPO can seamlessly extend to f-divergence, we also follow prior works to reparameterize our general exploratory bonus. The detailed algorithm can be formulated as in Algorithm 1.
Algorithm 1 Iterative Online Algorithm with Exploratory Bonus
0: Reference model $\pi_{\textrm{ref}}$, iteration number $T$, prompt set for each interation $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$, reward function $r$;
0: Trained model $\pi_{T}$;
1:for iteration t = 1, 2, . . . , T do
2:  for $x\in\mathcal{D}_{t}$ do
3:   $y_{1},y_{2}\sim\pi_{\textrm{ref}}(\cdot|x)$ and obtain the rewards $r(y_{1}),r(y_{2})$;
4:   Rank the reward and denote $y^{+},y^{-}$ as the preferred and dispreferred response between $y_{1},y_{2}$ and update $D_{t}=\{x,y^{w},y^{l}\}$;
5:   $\pi_{t}=\arg\min_{\pi}\mathcal{L}_{\textrm{DPO}}-\kappa\mathcal{L}_{bonus}(\pi)$
6:   update $\pi_{\textrm{ref}}$ with $\pi_{t}$ (optional)
7:  end for
8:end for
Hyperparameter settings and environments.
All experiments are conducted on two NVIDIA H200 GPUs. When training and sampling, the max length is set to 2048. For training, the batch size per device is set to 2; we enable the gradient checkpointing and the gradient accumulation step is set to 64; the learning rate is 5e-7 with a cosine scheduler, and the warm-up ratio is 0.03. In the main experiments, we use the best performance with $\kappa$ with a suitable ratio range to f-dpo loss across 1,1e2,1e4,1e6,1e81,1e-2,1e-4,1e-6,1e-8. For sampling, the temperature is set to 1. For in-domain evaluation and MATH evaluation, we set temperature to 0.6 and top-p to 0.9; we use the default setting of alpaca-eval.
Appendix E Proofs of Lemma 4.1
Lemma 4.1
Denote two objectives as $h(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}uf^{\prime}(u)-f(u)$ and $\hat{h}(u)=\mathbb{E}_{x\sim\rho,y\sim\pi_{\textrm{ref}}}\frac{u}{Z(x)}f^{\prime}(u)-f(\frac{u}{Z(x)})$, where $u$ varies with $\pi$ and $\pi_{\textrm{ref}}$. When $f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z(x)})]/[Z(x)uf^{\prime\prime}(u)]=\Lambda(x)$ is constant in $y$ and $\Lambda(x)>0$, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induce the same class of policies.
Proof  For a fixed $x$, with the Lagrange multiplier and KKT conditions, when $\pi(y|x)>0$, we obtain
$\frac{\partial h}{\partial\pi(y|x)}$ $=u(x,y)f^{\prime\prime}(u(x,y))\cdot\frac{\partial\pi_{\textrm{ref}}(y|x)u(x,y)}{\partial\pi(y|x)}=\mu_{1}(x),$ (31)
where $\mu_{1}(x)$ is a dual variable with respective to $x$. When $\frac{f^{\prime}(u)+uf^{\prime\prime}(u)-f(\frac{u}{Z})}{Zuf^{\prime\prime}(u)}=\Lambda(x)$, we have
$\mu_{1}(x)\Lambda(x)=\frac{1}{Z}(f^{\prime}(u)+uf^{\prime\prime}(u)-f^{\prime}(\frac{u}{Z}))\cdot\frac{\partial\pi_{\textrm{ref}}(y|x)u(x,y)}{\partial\pi(y|x)},$ (32)
which directly follows by the KKT conditions of $\hat{h}(x)$, i.e. $\frac{\partial\hat{h}}{\partial\pi(y|x)}=\mu_{2}(x)$ where $\mu_{2}(x)$ is a dual variable equals to $\mu_{1}(x)\Lambda(x)$. Hence, every policy that satisfies the stationary condition for $h$ also satisfies it for $\hat{h}$. Since $\Lambda(x)>0$, the second-order derivative $\frac{\partial h}{\partial^{2}\pi(y|x)}$ and $\frac{\partial\hat{h}}{\partial^{2}\pi(y|x)}$ have the same sign, which indicates they share the same local minima. Hence, minimizing the two objectives $\min_{\pi}-h(u)$ and $\min_{\pi}-\hat{h}(u)$ induces the same class of policies.
$\square$